Calculate the freezing and boiling point of a solution prepared by dissolving 32.60 g of ferric chloride (FeCl3) in 0.4 m3 of water (H2O).



1
Expert's answer
2022-05-09T17:03:04-0400

n(FeCl3) = 32.6 g / 162.2 g/mol = 0.2 mol

marity = 0.2 mol / 400 L = 0.0025 mol/L

freezing point= 0 - m*K = 0 - 0.0025*1.86 = -0.0047°С

boiling point= 100+m*E = 100 + 0.0025*0.52 = 100.0013°С